Solve the given problem, Be sure to show your complete solution.


  3Cu + 8HNO3 --> 3Cu(NO3)2 + 4H2O + 2NO


a) How many moles of copper(II) nitratrle can be prepared from 17.0 moles of Cu?

b) How many grams of copper(II) nitrate can be prepared using 3.8 moles of HNO3?

c) What mass of water results from the reaction of 8.50 kg of copper metal.


1
Expert's answer
2021-04-19T05:08:34-0400

We have given the reaction,

"3Cu + 8HNO_3 \rightarrow 3Cu(NO_3)_2 + 4H_2O + 2NO"

a.) 3 moles of "Cu" gives 3 moles of "Cu(NO_3)_2"

Hence, 17 moles of "Cu" gives 17 moles of "Cu(NO_3)_2"


b.) 8 moles of "HNO_3" gives 3 moles of "Cu(NO_3)_2"

3.8 moles of "HNO_3" gives "\dfrac{3}{8} \times 3.8" moles of "Cu(NO_3)_2" = 1.425 moles of "Cu(NO_3)_2"

Hence, 3.8 moles of "HNO_3" gives "1.425 \times 187.56" = "267.273g" of "Cu(NO_3)_2"


c.) 3 moles of Cu gives 4 moles of "H_2O"

133.76 moles of "Cu" produces "\dfrac{4}{3} \times 133.76" moles of "H_2O"

Hence, 133.76 moles of "Cu" produces "\dfrac{4}{3} \times 133.76 \times 18" = "3210.24g" of "H_2O"